OpenStudy (anonymous):

Which equation represents an asymptote of this hyperbola? ((x - 1)^2)/6^2 - ((y - 2)^2)/8^2 = 1

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

y = 2 + (4x)/3

OpenStudy (anonymous):

y = (2 + 4x)/3

OpenStudy (anonymous):

y = 2 + (3x)/4

OpenStudy (anonymous):

y = (2 + 3x)/4

OpenStudy (michele_laino):

hint: if we make this traslation: x-1=X y-2=Y where X, Y are the new coordinates, then we can rewrite your hyperbola as below: \[\frac{{{X^2}}}{{{6^2}}} - \frac{{{Y^2}}}{{{8^2}}} = 1\]

OpenStudy (michele_laino):

now, the asymptotes of that hyperbola are: \[Y = \frac{8}{6}X = \frac{4}{3}X\]

OpenStudy (anonymous):

Y = (4 x)/3

OpenStudy (anonymous):

so it would be Y=2+4x/3

OpenStudy (michele_laino):

so we have: \[y - 2 = \pm \frac{4}{3}\left( {x - 1} \right)\] I have used +/- since we have 2 asymptotes, namely: \[Y = \pm \frac{4}{3}X\]

OpenStudy (anonymous):

so it would be y = (2 + 4x)/3

OpenStudy (michele_laino):

yes! It is one of our asymptotes

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thanks!